Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by arvindsekar » Sun Jun 23, 2013 12:47 pm
What is the probability of getting exactly 2 sixes on three rolls of a fair six-sided die?

Can anyone explain how to solve this question?
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Sun Jun 23, 2013 1:33 pm
arvindsekar wrote:What is the probability of getting exactly 2 sixes on three rolls of a fair six-sided die?
There are 3 different ways to get exactly two 6's.
#1) 6 on 1st roll, 6 on 2nd roll, non-6 on 3rd roll
#2) 6 on 1st roll, non-6 on 2nd roll, 6 on 3rd roll
#3) non-6 on 1st roll, 6 on 2nd roll, 6 on 3rd roll

So, P(exactly two 6's) = P(6 on 1st, 6 on 2nd, non-6 on 3rd OR 6 on 1st, non-6 on 2nd, 6 on 3rd OR non-6 on 1st, 6 on 2nd, 6 on 3rd)
= P(6 on 1st, 6 on 2nd, non-6 on 3rd) + P(6 on 1st, non-6 on 2nd, 6 on 3rd) + P(non-6 on 1st, 6 on 2nd, 6 on 3rd)
= (1/6)(1/6)(5/6) + (1/6)(5/6)(1/6) + (5/6)(1/6)(1/6)
= 5/216 + 5/216 + 5/216
= 15/216
= [spoiler]5/72[/spoiler]



Aside: notice that P(6 on 1st, 6 on 2nd, non-6 on 3rd) = P(6 on 1st, non-6 on 2nd, 6 on 3rd) = P(non-6 on 1st, 6 on 2nd, 6 on 3rd)
So, once you find one of these probabilities, then you automatically know the other two.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by mgm » Mon Jun 24, 2013 4:36 am
Brent,

Thanks for the explanation , do you happen to have a list of good probability questions posted on btg handy ? For example in this question if the events were not independent then solution will be different. I really found it helpful to solve from your list of overlapping sets questions ...
Join the discussion

by Brent@GMATPrepNow » Mon Jun 24, 2013 6:07 am
If you want to focus on one topic at a time, you can use BTG's tagging feature to isolate one concept. For example, here are all of the questions tagged as probability questions: https://www.beatthegmat.com/forums/tags/ ... robability

See the left side of that linked page for more tag options.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Yaj » Wed Jul 10, 2013 11:59 am
Hi Brent,

Your explanation was really helpful!

I have a doubt (since I am just getting started, please excuse the v basic Qs I shoot) -

Why are we looking at OR (+) & not AND (*) here = 5/216 + 5/216 + 5/216 ? Since we are looking at all conditions working out (i.e 1/6*1/6*5/6) THREE times in a row, would not AND come in here ?

Thanks in advance
Join the discussion

by Brent@GMATPrepNow » Wed Jul 10, 2013 3:23 pm
Yaj wrote:Hi Brent,

Your explanation was really helpful!

I have a doubt (since I am just getting started, please excuse the v basic Qs I shoot) -

Why are we looking at OR (+) & not AND (*) here = 5/216 + 5/216 + 5/216 ? Since we are looking at all conditions working out (i.e 1/6*1/6*5/6) THREE times in a row, would not AND come in here ?

Thanks in advance
The solution requires ANDs and ORs. I highlighted the ORs because this is the part that students often miss.

As you suggested, parts such as P(6 on 1st, 6 on 2nd, non-6 on 3rd) are actually AND probabilities.
So, P(6 on 1st, 6 on 2nd, non-6 on 3rd) = P(6 on 1st AND 6 on 2nd AND non-6 on 3rd)
= P(6 on 1st) x P(6 on 2nd) x P(non-6 on 3rd)
= 1/6 x 1/6 x 6/6
= 5/216

I hope that helps.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by send2dar » Sat Jul 13, 2013 6:25 am
Brent, I liked your explanation...
Thanks
---------------
V for Victory, V for Verbal. Only performance in Verbal section separates me from the Victory I want (Trying for 750+)
Join the discussion

by Matt@VeritasPrep » Sat Jul 13, 2013 9:56 am
Just to add one more point to this discussion:

Once you figure out the probability of getting exactly two sixes in some order - say a 6, a 6, then a non-6 - all you have to do is figure out how the number of possible arrangements, then multiply the probability of one arrangement by the number of arrangements.

For example, the odds of going

Six, Six, Non-Six

are 1/6 * 1/6 * 5/6

There are three ways to arrange the letters SSN, so there are three possible arrangements.

So the probability is 1/6 * 1/6 * 5/6 * 3.

The more complex these sort of problems get, the better it is to think of the arrangements combinatorially.
Join the discussion

by vipulgoyal » Sun Jul 14, 2013 11:03 pm
experts I was trying to get it by (1 - odds = desired) approach, tried multiple times, please shed some light, 1 - {p(getthing one 6) + P(getting 3 6)}
Join the discussion

by Brent@GMATPrepNow » Mon Jul 15, 2013 6:21 am
vipulgoyal wrote:experts I was trying to get it by (1 - odds = desired) approach, tried multiple times, please shed some light, 1 - {p(getting one 6) + P(getting 3 6)}
Yeesh - that's a lot of work. But let's try it out to show that both ways do, indeed, work:

P(exactly two 6's) = 1 - P(not two 6's)
= 1 - P(zero 6's OR one 6 OR three 6's)
Aside: You forgot to mention the option of zero 6's in your post

Let's calculate P(zero 6's OR one 6 OR three 6's)
P(zero 6's OR one 6 OR three 6's) = P(zero 6's) + P(one 6) + P(three 6's)
= (5/6)(5/6)(5/6) + {(1/6)(5/6)(5/6) + (5/6)(1/6)(5/6) + (5/6)(5/6)(1/6)} + (1/6)(1/6)(1/6)
= 125/216 + {75/216} + 1/6
= 201/216

So, P(exactly two 6's) = 1 - 201/216
= 15/216
= [spoiler]5/72[/spoiler]

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion